# HILCodec: High Fidelity and Lightweight Neural Audio Codec

### v1

@@ III-B2 Variance-Constrained Residual Block
In this subsection, we present the Variance-Constrained Residual Block (VCRB) designed to downscale the variance of the output of each residual branch such that the variance increases linearly with the network depth. It is motivated by the NFNet block [62] in image domain. A VCRB can be mathematically written as
@@
where $\alpha$ and $\beta_{n}$ are fixed scalars to be discussed. We consider the residual branch $f_{n}$ right after initialization. Also, we assume that the network input $\mathbf{x}_{0}$ consists of i.i.d. random variables with a mean of 00 and a standard deviation of $1$, and $f_{n}(\cdot)$ is a sample function of a random process indexed by $\mathbf{x}$. There are three key components in (2). Firstly, we randomly initialize the parameters of $f_{n}$ such that $\mathbf{Var}(f_{n}(\mathbf{x}_{n}))=\mathbf{Var}(\mathbf{x}_{n})$. More specifically, when a convolution layer follows an activation function, we apply He initialization [48]. When it doesn’t, we apply LeCun initialization [49] (see Appendix A). Secondly, we analytically calculate $\mathbf{Var}(\mathbf{x}_{n})$ (see (6)) and set
@@
This ensures that the variance of the input (and also the output) of $f_{n}(\cdot)$ is $1$:
@@
III-B3 Other Building Blocks
For the spectrogram blocks described in Section III-A4, we treat them as a special type of residual block. This means that similar to the other residual blocks, we scale the output of each spectrogram block $g(\cdot)$ as:
@@ III-B3 Other Building Blocks
III-B4 Input/Output Normalization
We further normalize the input and output of both the encoder and decoder. We compute the mean and standard deviation of $10000$ randomly sampled audio chunks from the training dataset. At the beginning of the encoder, we insert a normalization layer using pre-calculated statistics. We also insert normalization layers for multi-resolution spectrogram inputs. In the decoder, right after the final convolution and prior to the hyperbolic tangent activation, we insert a de-normalization layer. The final element we normalize is the output of the encoder (equivalently the input of the decoder before quantization), denoted as $\mathbf{z}$. Since we apply $l_{2}$-normalization before quantization as discussed in Section III-A3, the mean square of $\mathbf{z}$ is $\mathbb{E}(\mathbf{z}^{2})=1/C$, where $C$ is the channel dimension of $\mathbf{z}$. Ignoring the square of the mean of $\mathbf{z}$, we multiply $\sqrt{C}$ to $\mathbf{z}$ after $l_{2}$-normalization so that the variance is approximately $1$ as given by
@@ III-B4 Input/Output Normalization
III-B5 Zero-Initialized Residual Branch
For deep residual networks, it has been empirically shown that initializing every residual block to be an identity function at the beginning of training can further enhance the performance [53]. Following [62], we add a scalar parameter initialized to zero at the end of each residual branch, which we also found beneficial for the fast convergence of the model. While this zero-initialization scheme results in a modification of $\beta_{n}$ in Section III-B2, we found that ignoring this effect does not detrimentally impact the performance of our model.
@@ IV-A1 Problems of previous discriminators
It is well-known that audio encompasses a significantly large number of samples within a short time span. For instance, an audio clip that is 3 seconds long at a sampling rate of 24kHz contains 72,000 samples. Consequently, waveform domain discriminators typically accept audio inputs after downsampling. As shown in Fig. 3(a) and (b), the Multi-Scale Discriminator (MSD) [22] uses average pooling, and the Multi-Period Discriminator (MPD) [24] relies on a plain downsampling. The combination of MSD and MPD has achieved significant success in the field of neural vocoders and has been employed in various neural audio codecs [41, 40, 42]. However, from a signal processing perspective, these discriminators induce distortion to the input signal. MPD relies on a plain downsampling, which causes aliasing of the signal. In the case of MSD, average pooling can be interpreted as a low-pass filter using a rectangular window. Fig. 3(d) depicts the magnitude response of a 4-tap average pooling used in MSD. The passband is not flat, the transition band is too wide and not steep, and the attenuation level in the stopband is too low. It is evident that average pooling distorts input signal and is insufficient to prevent aliasing caused by subsequent downsampling. Furthermore, after average pooling, MSD is limited to modeling low-frequency components only. This highlights the need for a more sophisticated waveform-domain discriminator capable of preventing distortion and modeling the entire frequency spectrum.
@@ IV-A2 Proposed discriminator
We propose a novel Multi-Filter Bank Discriminator (MFBD). As illustrated in Fig. 3(c), it initially transforms the waveform into $N$ sub-bands using a filter bank, followed by a downsampling with a factor of $N$. Each sub-band is then fed into a single discriminator with shared weights. This enables the discriminator to observe the entire frequency components with minimum distortion. Similar to [24], MFBD is composed of multiple sub-discriminators with the number of sub-bands $N\in\{1,2,3,5,7,11\}$. This approach has two beneficial effects. Sub-discriminators with small $N$ can model short-term dependencies, while those with large $N$ can model long-term dependencies. Additionally, because we use a coprime number of sub-bands, even the distortion in transition bands can be covered by other sub-discriminators. The model architecture of MFBD is configured in the same way as MPD.
Similar approach was also introduced in Avocodo [27]. It uses pseudo-quadratic mirror filter bank [7] to obtain sub-band features and feed the entire features to a discriminator. The key differences are that ours have more generalization ability by using a shared discriminator across different sub-bands. We also utilize multiple sub-discriminators with co-prime number of sub-bands to cope with distortion in transition bands. Our experiments demonstrate the superiority of MFBD over discriminators of Avocodo (see Section VI-B). For more detailed comparison, please refer to Appendix B.
@@ IV-B Multi-Resolution Spectrogram Discriminator
Along with MFBD, we also apply the multi-resolution spectrogram discriminator (MRSD). Simultaneously using both waveform domain discriminator and time-frequency domain discriminator has been proven to enhance the audio quality [38, 40]. We obtain complex spectrograms with Fourier transform sizes of $\{128,256,512,1024\}$ and hop lengths of $\{32,64,128,256\}$. We represent each spectrogram as a real-valued tensor with two channels (real and imaginary) and pass it through a discriminator composed of 2D convolutions.
@@ V-A4 Commitment Loss
$\mathcal{L}_{c}=\left\|\mathcal{G}_{e}(x)-\mathcal{Q}\Big{(}\text{sg}[\mathcal{G}_{e}(x)]\Big{)}\right\|_{2}^{2},$ (13)
@@ V-A5 Loss Balancer
~$\widetilde{\nabla}_{w}=\left(\sum_{i}\lambda_{i}\cdot\frac{g_{i}}{\mathbb{E}\left[\|g_{i}\|_{2}\right]}\right)\frac{\partial\hat{x}}{\partial w}+\lambda_{c}\cdot\frac{\partial\mathcal{L}_{c}}{\partial w},$ (16)
@@ V-B Datasets and Training Settings
We trained two versions of HILCodec: one trained on clean speech, and the other trained on general audio, including clean speech, noisy speech and music. For clean speech datasets, we used speech segments from DNS Challenge 4 [63] and the VCTK-0.92 corpus [64]. For noisy speech, we mixed clean speech segments with noise segments from DNS Challenge 4 during training on-the-fly. For music, we used Jamendo [65]. All audio files were downsampled to a sampling rate of 24kHz. During training on general audio, we sampled each type of audio with the same probability. We randomly extracted 1 second long segment from each audio file, applied a random gain in [-10dB, 6dB], and scaled down the audio when a clipping occurred during a mixing or a random gain adjustment.
@@ V-C Baselines
We compare HILCodec to various end-to-end neural audio codecs that are publicly available. EnCodec [39] and Descript Audio Codec [42] are models trained on general audio. HiFi-Codec [41] and AudioDec [40] are models trained on clean speech. For all models, we used versions with the same sampling rate of 24kHz. We also included EVS [9], the signal processing-based audio codec which has demonstrated state-of-the-art performance among traditional codecs [10]. It is a hybrid codec that uses LPC when compressing speech and uses transform coding when compressing the others. Since it does not support the sampling rate of 24kHz, we upsampled the audio to 32kHz, compressed with EVS, and downsampled back to 24kHz.
@@ V-D1 Subjective Metric
To compare HILCodec with the baselines, we conducted MUSHRA test [66] using webMUSHRA framework [68]. As reference audios, we selected four clean speeches, four noisy speeches, and four music pieces for general audio coding, and 12 clean speeches for speech coding. For each test item, we included a low anchor, which was a low-pass filtered version of the reference audio with a cut-off frequency of 3.5kHz, but excluded a mid anchor with a cut-off frequency of 7kHz. Post-screening was performed to exclude assessors who rated hidden references lower than a score of 90 more than 15% of the test items or rated low anchors higher than 90 more than 15% of the test items. The number of assessors after post-screening was 11.
@@ V-D2 Objective Metric
For model developments and ablation studies, we utilized ViSQOL Audio v3 [67]. We selected 100 audio samples for each of three categories, upsampled reference audios and coded audios to a sampling rate of 48kHz, and calculated ViSQOL scores.
@@ VI-A Subjective Quality Comparison
Fig. 4 illustrates the MUSHRA scores obtained from HILCodec and the other baseline codecs. Note that HILCodec, EnCodec, AudioDec, and EVS are streamable while Descript and HiFi-Codec are not. Also, HILCodec has a complexity smaller than most of the other neural codecs (see Section VI-C). In Fig. 4(a), we can see that for general audio coding, HILCodec at 3kbps significantly outperforms Descript at 3kbps, EnCodec at 3kbps, and EVS at 5.9kbps. Moreover, in Fig. 4(b), HILCodec at 9kbps surpasses other neural audio codecs at 9kbps and EVS at 16.4kbps. In Fig. 4(c), for clean speech coding, HILCodec at 3kbps has slightly lower sound quality than high-complexity unstreamable HiFi-Codec at 3kbps. However, it outperforms AudioDec at 6.4kbps and EVS at 5.9kbps by significant margins.
@@ VI-B Ablation Studies
TABLE I: Ablation studies for the generator. Mean and 95% confidence interval values are reported.
Ablation ViSQOL
@@
We performed ablation studies to prove the effectiveness of the proposed techniques. Because of training resource limits, all models presented in this subsection were trained with a smaller timestep of 156k iterations.
VI-B1 Generator
Starting from HILCodec trained on general audio, we removed each generator component while maintaining the other components and measured ViSQOL scores. The results are presented in Table I. Regarding the architectural choice, removing $L_{2}$-normalization and spectrogram blocks each results in quality degradation. Moreover, replacing the variance-constrained residual blocks with plain residual blocks, removing input/output normalization, and removing zero initialization each results in a lower ViSQOL score. This indicates that each component we propose has a beneficial effect on the audio quality.
@@ VI-B2 Discriminator
TABLE II: MUSHRA scores of different discriminators. Mean and 95% confidence interval values are reported.
@@
MPD+MSD (HiFi-GAN) 69.32 $\pm$ 2.95
CoMBD+SBD (Avocodo) 74.66 $\pm$ 3.34
MFBD+MRSD (HILCodec) 82.43 $\pm$ 2.27
We compared the subjective quality of the proposed discriminators with that of MPD and MSD used in HiFi-GAN, and CoMBD and SBD used in Avocodo. For MPD and MSD, we used the same generator and training configurations to ensure a fair comparison. However, since CoMBD operates on multiple resolutions, we modified the generator so that the last three decoder blocks generate waveforms with different sampling rates, as in Avocodo. This made it impossible to use the loss balancer. Instead, we strictly followed the loss functions and loss coefficients used in [27]. Table II shows the MUSHRA scores for different combinations of discriminators. MFBD and MRSD provide better perceptual quality than the other discriminators.

### v2

@@ III-E2 Variance-Constrained Residual Block
In this subsection, we present the Variance-Constrained Residual Block (VCRB) designed to downscale the variance of the output of each residual branch such that the variance increases linearly with the network depth right after initialization. It is motivated by the NFNet block [62] in image domain. A VCRB can be mathematically written as
@@
where $\alpha$ and $\beta_{n}$ are fixed scalars to be discussed. We consider the residual branch $f_{n}$ right after initialization. Also, we assume that the network input $\mathbf{x}_{0}$ consists of i.i.d. random variables with a mean of 00 and a standard deviation of $1$, and $f_{n}(\cdot)$ is a sample function of a random process indexed by $\mathbf{x}$. There are three key components in (2). Firstly, we carefully initialize the parameters of $f_{n}$ such that $\mathbf{Var}(f_{n}(\mathbf{x}_{n}))=\mathbf{Var}(\mathbf{x}_{n})$ (details in Appendix C). Secondly, we analytically calculate $\mathbf{Var}(\mathbf{x}_{n})$ as in (6), and set
@@
This ensures that the variance of the input and the output of $f_{n}(\cdot)$ is $1$:
@@
III-E3 Other Building Blocks
For the spectrogram blocks described in Section III-D, we treat them as a special type of residual block. This means that similar to the other residual blocks, we scale the output of each spectrogram block $g(\cdot)$ as:
@@ III-E3 Other Building Blocks
III-E4 Input/Output Normalization
We further normalize the input and output of both the encoder and decoder. We compute the mean and standard deviation of $10000$ randomly sampled audio chunks from the training dataset. At the beginning of the encoder, we insert a normalization layer using pre-calculated statistics. We also insert normalization layers for multi-resolution spectrogram inputs. In the decoder, right after the final convolution and prior to the hyperbolic tangent activation, we insert a de-normalization layer. The final element we normalize is the output of the encoder (equivalently the input of the decoder before quantization), denoted as $\mathbf{z}$. Since we apply $l_{2}$-normalization before quantization as discussed in Section III-C, the mean square of $\mathbf{z}$ is $\mathbb{E}(\mathbf{z}^{2})=1/C$, where $C$ is the channel dimension of $\mathbf{z}$. Ignoring the square of the mean of $\mathbf{z}$, we multiply $\sqrt{C}$ to $\mathbf{z}$ after $l_{2}$-normalization so that the variance is approximately $1$ as given by
@@ III-E4 Input/Output Normalization
III-E5 Zero-Initialized Residual Branch
For deep residual networks, it has been empirically shown that initializing every residual block to be an identity function at the beginning of training can further enhance the performance [53, 72]. Following [62, 71], we add a scalar parameter initialized to zero at the end of each residual branch, which we also found beneficial for the fast convergence of the model. While this zero-initialization scheme results in a modification of $\beta_{n}$ in Section III-E2, we found that ignoring this effect does not detrimentally impact the performance of our model.
@@ IV-A1 Problems of previous discriminators
Digital audio encompasses a significantly large number of samples within a short time span. For instance, an audio clip that is 3 seconds long at a sampling rate of 24kHz contains 72,000 samples. Consequently, waveform domain discriminators typically accept audio inputs after downsampling. As shown in Fig. 3(a) and (b), the Multi-Scale Discriminator (MSD) [22] uses average pooling, and the Multi-Period Discriminator (MPD) [24] relies on a plain downsampling. The combination of MSD and MPD has achieved significant success in the field of neural vocoders and has been employed in various neural audio codecs [41, 40, 42]. However, from a signal processing perspective, these discriminators induce distortion to the input signal. MPD relies on a plain downsampling, which causes aliasing of the signal. In the case of MSD, average pooling can be interpreted as a low-pass filter using a rectangular window. Fig. 3(d) depicts the magnitude response of a 4-tap average pooling used in MSD. The passband is not flat, the transition band is too wide and not steep, and the attenuation level in the stopband is too low. It is evident that average pooling distorts input signal and is insufficient to prevent aliasing caused by subsequent downsampling. Furthermore, after average pooling, MSD is limited to modeling low-frequency components only. This highlights the need for a more sophisticated waveform domain discriminator capable of preventing distortion and modeling the entire frequency spectrum.
@@ IV-A2 Proposed discriminator
We propose a novel Multi-Filter Bank Discriminator (MFBD). As illustrated in Fig. 3(c), it initially transforms the waveform into $N$ sub-bands using a filter bank, followed by a downsampling with a factor of $N$. Each sub-band is then fed into a single discriminator with shared weights. This enables the discriminator to observe the entire frequency components with minimum distortion. We used pseudo-quadratic mirror filters (PQMF) [7] as the filter bank. Similar to [24], MFBD is composed of multiple sub-discriminators with the number of sub-bands $N\in\{1,2,3,5,7,11\}$. This approach has two beneficial effects. Sub-discriminators with small $N$ can model short-term dependencies, while those with large $N$ can model long-term dependencies. Additionally, because we use a coprime number of sub-bands, even the distortion in transition bands can be covered by other sub-discriminators. The model architecture of MFBD is configured in the same way as MPD.
A similar approach was also introduced in Avocodo [27]. It uses PQMF to obtain sub-band features and feeds the entire features to a discriminator. The key differences are that MFBD uses a shared discriminator across different sub-bands, while Avocodo uses distinct discriminators for each sub-band. We also utilize multiple sub-discriminators with a co-prime number of sub-bands to cope with distortion in transition bands. Our experiments demonstrate the superiority of MFBD over the discriminators used in Avocodo (Section VI-B2). For more detailed comparisons, please refer to Appendix D.
@@ IV-B Multi-Resolution Spectrogram Discriminator
Along with MFBD, we also apply the multi-resolution spectrogram discriminator (MRSD). Simultaneously using both waveform domain discriminator and time-frequency domain discriminator has been proven to enhance the audio quality [38, 40]. We obtain complex spectrograms with Fourier transform sizes of $\{128,256,512,1024\}$ and hop lengths of $\{32,64,128,256\}$. We represent each spectrogram as a real-valued tensor with two channels (real and imaginary) and pass it through a discriminator composed of six 2D convolutions. The model architecture is almost the same as that of [39], except that ours has output channels of $\{16,16,32,64,128,1\}$ instead of $\{32,32,32,32,32,1\}$ used in [39]. We found that this modification results in faster training while maintaining similar audio quality.
@@ V-A4 Commitment Loss
$\mathcal{L}_{c}=\left\|\mathcal{G}_{e}(x)-\mathcal{Q}\Big{(}\text{sg}\big{[}\mathcal{G}_{e}(x)\big{]}\Big{)}\right\|_{2}^{2},$ (13)
@@ V-A5 Loss Balancer
~$\widetilde{\nabla}_{w}=\left(\sum_{i}\lambda_{i}\cdot\frac{g_{i}}{\mathbb{E}\big{[}\|g_{i}\|_{2}\big{]}}\right)\frac{\partial\hat{x}}{\partial w}+\lambda_{c}\cdot\frac{\partial\mathcal{L}_{c}}{\partial w},$ (16)
@@ V-B Datasets and Training Settings
We trained two versions of HILCodec: one trained on clean speech, and the other trained on general audio, including clean speech, noisy speech and music. For clean speech datasets, we used speech segments from DNS Challenge 4 [63] and the VCTK-0.92 corpus [64]. For noisy speech, we mixed clean speech segments with noise segments from DNS Challenge 4 during training on-the-fly. For music, we used Jamendo [^2] It is distributed in 44.1kHz MP3 format. MP3 compression often discards high-frequency components. However, we downsampled the training dataset to 24kHz. Therefore, we suspect that using the training dataset in MP3 format is not problematic. [65]. All audio files were downsampled to a sampling rate of 24kHz. The total amount of training data is 2,548 hours for clean speech, 175 hours for noise, and 3,730 hours for music. During training on general audio, we sampled each type of audio with the same probability. We randomly extracted a one-second-long segment from each audio file, applied a random gain in the range of [-10dB, 6dB], and scaled down the audio when a clipping occurred during a mixing or a random gain adjustment.
@@ V-C Baselines
We compare HILCodec to various end-to-end neural audio codecs that are publicly available. EnCodec [39] and Descript Audio Codec [42] are models trained on general audio. HiFi-Codec [41] and AudioDec [40] are models trained on clean speech. For all models, we used versions with a sampling rate of 24kHz. We also included EVS [9], the signal processing-based audio codec which has demonstrated state-of-the-art performance among traditional codecs [10]. It is a hybrid codec that uses LPC when compressing speech and uses transform coding when compressing the others. Since it does not support the sampling rate of 24kHz, we upsampled input audios to 32kHz, compressed with EVS, and downsampled back to 24kHz.
@@ V-D1 Subjective Metric
To compare HILCodec with the baselines, we conducted MUSHRA listening tests [66] using webMUSHRA framework [68]. As reference audios, we selected four clean speeches, four noisy speeches, and four music pieces for general audio coding, and 12 clean speeches for speech coding. For each test item, we included a low anchor, which was a low-pass filtered version of the reference audio with a cut-off frequency of 3.5kHz, but excluded a mid anchor with a cut-off frequency of 7kHz. Post-screening was performed to exclude assessors who rated hidden references lower than a score of 90 more than 15% of the test items or rated low anchors higher than 90 more than 15% of the test items. The number of assessors after post-screening was 11.
@@ V-D2 Objective Metric
Conducting listening tests is the most reliable way to measure audio quality, but it is time-consuming and costly. Therefore, for model developments and ablation studies, we utilized ViSQOL Audio v3 [67]. We selected 100 audio files at a sampling rate of 24kHz for each of the three categories, upsampled reference and coded audio files to a sampling rate of 48kHz, and calculated ViSQOL scores. Further objective evaluations are provided in Appendix A.
@@ VI Results
Figure 5: MUSHRA scores for different audio types with 95% confidence intervals. (d) ”w/o Arch” denotes HILCodec without $L_{2}$-normalization and spectrogram blocks, ”w/o VCD” denotes HILCodec without the variance-constrained design, and ”w/o VCD, w/ BN” denotes HILCodec without the variance-constrained design and with batch normalization layers.
@@ VI-A Subjective Quality Comparison
Fig. 4 illustrates the MUSHRA scores obtained from HILCodec and the other baseline codecs. Note that HILCodec, EnCodec, AudioDec, and EVS are streamable while Descript and HiFi-Codec are not. Also, HILCodec has a complexity smaller than most of the other neural codecs (Section VI-C). In Fig. 4(a), for general audio coding, HILCodec at 3kbps significantly outperforms Descript at 3kbps, EnCodec at 3kbps, and EVS at 5.9kbps. Moreover, in Fig. 4(b), HILCodec at 9kbps surpasses other neural audio codecs at 9kbps and EVS at 16.4kbps. In Fig. 4(c), for clean speech coding, HILCodec at 3kbps has slightly lower sound quality than high-complexity unstreamable HiFi-Codec at 3kbps. However, it outperforms AudioDec at 6.4kbps and EVS at 5.9kbps by significant margins. Fig. 5(a) and (b) show MUSHRA scores for different audio types. HILCodec outperforms other codecs across all bitrates and audio types.
@@ VI-B Ablation Studies
We conducted ablation studies to prove the effectiveness of the proposed techniques. Due to training resource limits, all models presented in this subsection were trained with a smaller timestep of 156k iterations. All models were trained on a general audio corpus and evaluated at 3kbps.
VI-B1 Generator
TABLE I: ViSQOL scores for ablation studies of the generator. Mean and 95% confidence interval values are reported.
Generator ablation ViSQOL
@@ VI-B1 Generator
TABLE II: MUSHRA scores for ablation studies of the generator. Mean and 95% confidence interval values are reported.
Generator ablation MUSHRA
HILCodec 80.58 $\pm$ 1.89
Without architectural changes 81.48 $\pm$ 2.32
Without VCD 77.15 $\pm$ 2.35
Without VCD, with BatchNorm 66.71 $\pm$ 2.44
@@
Starting from HILCodec, we removed each proposed generator component while maintaining the other components and measured ViSQOL scores. The results are presented in Table I. Regarding the architectural changes, removing $L_{2}$-normalization and spectrogram blocks each results in quality degradation. Moreover, replacing the variance-constrained residual blocks with plain residual blocks, removing input/output normalization, and removing zero initialization each results in a lower ViSQOL score.
We conducted a MUSHRA test for the ablation study. Comparing too many models at once increases the difficulty of the MUSHRA test and the burden on assessors. Therefore, we compared four models as follows - HILCodec, HILCodec without the proposed architectural changes, HILCodec without VCD, and HILCodec without VCD and with Batch Normalization. The MUSHRA scores are presented in Table II, and the MUSHRA scores for different audio types are presented in Fig. 5(c). It is important to note that mixed-precision training failed for all models except the baseline HILCodec. We made three observations. Firstly, in Table II, the difference in MUSHRA scores between HILCodec and HILCodec without architectural changes is within the margin of error. Upon closer examination, the proposed architectural changes degrade the clean speech quality while improving the music quality. We suspect this is a characteristic of $L_{2}$-normalization. Practitioners may choose whether to use $L_{2}$-normalization depending on their audio type of interest. However, one should keep in mind that discarding $L_{2}$-normalization resulted in the failure of mixed-precision training. Secondly, when VCD is removed, the MUSHRA score degrades significantly. The audio quality remains similar for clean speech, slightly degrades for noisy speech, and significantly degrades for music. This validates the beneficial effect of VCD on the audio quality. Lastly, using Batch Normalization instead of VCD definitely degrades the subjective quality. This implies that Batch Normalization is ill-suited for the SEANet architecture, providing the motivation for VCD (Section III-E1).
@@ VI-B2 Discriminator
TABLE III: MUSHRA scores of different discriminators. Mean and 95% confidence interval values are reported.
@@
MFBD+MRSD (HILCodec) 82.81 $\pm$ 1.96
MPD+MRSD 72.72 $\pm$ 2.84
MSD+STFTD (SoundStream) 75.82 $\pm$ 2.50
CoMBD+SBD (Avocodo) $7$$\pm$ 2.62
MPD+MSD (HiFi-GAN) 73.47 $\pm$ 2.62
To demonstrate the effectiveness of the proposed discriminators (MFBD and MRSD), we compared the subjective qualities of various discriminators while maintaining the same generator and training configurations. We included various combinations of discriminators from previously successful neural vocoders and neural audio codecs: MPD and MSD from HiFi-GAN, CoMBD and SBD from Avocodo, and MSD and a STFT discriminator (STFTD) from SoundStream. We also included the combination of MPD and MRSD to validate the superiority of MFBD. Since CoMBD operates on multiple resolutions, we modified the generator for the combination of CoMBD and SBD so that the last three decoder blocks generate waveforms with different sampling rates, as in Avocodo. This modification made it impossible to use the loss balancer. Instead, we strictly followed the loss functions and loss coefficients used in [27].
Table III presents the MUSHRA scores for different combinations of discriminators. The combination of MFBD and MRSD offers the best perceptual quality. Moreover, replacing the proposed MFBD with MPD reduces the subjective quality. Given that the only difference between MFBD and MPD is the use of a filter bank before downsampling, this result demonstrates that our simple solution leads to a considerable improvement in audio quality.
Fig. 5(d) illustrates the MUSHRA scores for different audio types. For clean speech, our discriminators and Avocodo’s discriminators both show the best quality, and the difference in scores between the two is within the margin of error. For noisy speech, our disrciminators show the best quality. For music, our discriminators outperform the others by large margins. These results demonstrate that the proposed discriminators are effective across various audio types.